Find the length of the portion of the parabola x=3t^2, y=6t cut off by the line 3x+y-3=0
1
Expert's answer
2020-03-25T12:38:10-0400

ANSWER : "\frac { \sqrt { 10 } }{ 3 } +3\sqrt { 2 } +3\ln { \frac { 1+\sqrt { 10 } }{ 3(\sqrt { 2 } -1) } }"

EXPLANATION.

The length L of the curve "G= \left\{ (x(t),y(t)):a\le x\le b \right\}" is calculated by the formula

"L=\int _{ a }^{ b }{ \sqrt { { \left[ x'(t) \right] }^{ 2 }+{ \left[ y'(t) \right] }^{ 2 } } } dt" . To determine the values of "a" and "b" we find the value of the parameter "t" corresponding to the coordinates of the points of intersection of the parabola and the line. To do this , we solve the equation "3(3t^2) +6t-3=0" or "(3t+1)^2-4=0" . Hence "a=-1, b=\frac { 1 }{ 3 }" ."x'(t)=6t,\quad y'(t)=6,\ \sqrt { { \left[ x'(t) \right] }^{ 2 }+{ \left[ y'(t) \right] }^{ 2 } } =\sqrt { 36{ t }^{ 2 }+36 }"

"L=6\int _{ -1 }^{ 1/3 }{ \sqrt { 1+t^{ 2 } } } dt" "=F(1\/3)-F(-1)" , where "F(t)= 6 \left [ {\frac{t}{2}} \sqrt{1+t^2} +\frac{1}{2} \ln \left | t+\sqrt{1+t^2} \right | \right ]" , "F(1/3)=\left[ \frac { \sqrt { 10 } }{ 3 } +3\ln { \frac { 1+\sqrt { 10 } }{ 3 } } \right]" ,"F(-1)=\left[ -3\sqrt { 2 } +3\ln { (\sqrt { 2 } -1) } \right]"